solve 6x-3y=5
x+2y=0

Answer by mananth(16946)

6x-3y=5.............1
x+2y=0 .............2

multiply (1)by 2
Multiply (2) by 3
12x-6y=10
3x+6y=0

Add the two equations 15
x=10

/15

x= 2/3

plug value of x In (1)

6x-3y= 5
4-3y= 5
-3y=5-4
-3y=1
y=- 1/3